Question 1.

For a hypothetical chemical reaction:

2 A ( g ) + 2 B ( g ) ⟶ p r o d u c t s

the rate law is first order with respect to [A] and second order with respect to [B]. If the rate of this reaction at a specific temperature was 4.1 x 10-2 M/s when the [A] = 1.2 M and [B] = 1.5 x 10-2 M, what is the rate constant (k) with the correct units?

Group of answer choices

A.) 1.5 x 102 M-2 s-1

B.) 1.0 x 102 M-1 s-1

C.) 1.5 x 102 s-1

D.)1.0 x 102 M-2 s-1

Question 2

The reaction:

X ⟶ Y

follows first-order kinetics. A certain amount of X is placed in a container and proceeds at a rate specific for the reaction and temperature. What would happen to the relative rate if the same amount of X was used but the volume of the container was reduced by half?

Group of answer choices

A.) The rate would be halved.

B.) The rate would quadruple.

C.) The rate would double.

D.)The rate would not change.
Answer #1

2 A ( g ) + 2 B ( g ) ⟶ p r o d u c t s

Rate = K[A]^1[B]^2

4.1*10^-2 = K(1.2)(1.5*10^-2)^2

K = 4.1*10^-2/(1.2)(1.5*10^-2)^2

K = 152M^-2s^-1

A. 1.5*10^2M^-2 s^-1 >>>>answer

C.) The rate would double. >>>answer

Rate = K[A]^1

Rate = K(no of mole/V)

v = 1/2

Rate is double